# Pierre Cartier: A Visionary Mathematician

### v1

@@ 1 Hopf algebras, group theory
Cartier had a guiding thread in his work that revolved around groups and Hopf algebras. He viewed Hopf algebras as a powerful tool for understanding various algebraic phenomena. His research often focused on their applications in algebraic topology, representation theory, and mathematical physics. The duality in Hopf algebras, where every algebraic structure has a corresponding co-algebra structure, resonated with Cartier’s broader mathematical philosophy, which emphasized the interconnectedness of different fields of mathematics.
@@ 1.1 On the theory of Formal groups
The essential progress was the introduction of curve modules associated with formal groups and the notion of a typical curve. This novel concept retrospectively explained the arithmetic properties of the classical exponential series and, in particular, the existence of the Artin-Hasse exponential. It also offered a flexible interpretation of the Dieudonné module and allowed me to refine known theorems on Witt groups. The universal typical curve in a Witt group plays a central role in these demonstrations.
@@ 1.2 Quantum groups
At that time, I was nearing the end of my mathematical research program and was in a period of waiting. I realized that quantum groups were something entirely new that could be approached with techniques I had previously used (like Lie groups and differential geometry), which I knew well. There was also motivation from physics, even if not always apparent, which reoriented my interests for a good ten years. In fact, I still organize a seminar at the École Polytechnique called ”Quantum Groups and Poisson Geometry,” which continues this work, twenty years later.
@@ 1.3 The Cosmic Galois Groups
The motivic Galois group deals with the automorphisms of certain transcendental numbers, which are similar to the constants appearing in Feynman diagram computations. This observation led me to interpret the group in [12] as a symmetry group governing the fundamental constants of physics. In the Standard Model, these constants are often adjusted empirically, without much mathematical explanation. I believe that this group could express new symmetries among these constants, which might have significant implications for cosmology. My ultimate dream is to unite the ideas in [12] with the motivic Galois group, even if, for now, it’s still an ongoing research program888personal communication [11].
@@
It took a lot of work and time to obtain a concrete realization of Pierre Cartier’s idea of the Cosmic Galois Group as well as its action on the coupling constants of renormalizable physical theories. The first step, following the introduction of the Hopf algebra of Feynman graphs, was the realization that the renormalization process is, in fact (in the dimensional regularization scheme), identical to the Birkhoff decomposition in pure mathematics, which occurs in the classification of vector bundles on the sphere (see [23]). Once this result was obtained, it again took time, using the Riemann-Hilbert correspondence, to realize ([24]) that there exists a universal Hopf algebra $\mathcal{H}$ behind all these computations999introducing the category of “equisingular flat connections,” which is shown to be a Tannakian category, meaning it is equivalent to the category of modules over a certain pro-algebraic group. This is the graded dual $\mathcal{H}$ of the universal enveloping algebra of the free graded Lie algebra generated by elements of degree $n$ for each positive $n$. Let $\mathbb{U}$ be the affine group scheme associated with $\mathcal{H}$, and let $\mathcal{H}$ and $\mathbb{U}^{*}=\mathbb{U}\rtimes\mathbb{G}_{m}$ be its semidirect product by the action of the multiplicative group given by the grading. It follows (see [24], Corollary 1.107) that this group acts on the coupling constants of any renormalizable physical theory, thus providing the best model for the Cosmic Galois Group. It gives a concrete realization of the symmetry group of physical theories and has a tantalizing similarity with the motivic Galois group of mixed Tate motives (see [24], Corollary 1.111).
@@ 2 His many key ideas offered to others
What attracts me to the epistemology of mathematics is understanding how mathematics is woven into civilization. I’m interested in how mathematical concepts emerge from societal concerns, reflecting the zeitgeist. One of my students specialized in the history of mathematics in China and worked on her thesis with me and my sinologist brother. André Weil sparked my interest in the history of mathematics, teaching me to view past mathematicians – like Euclid, Archimedes, Fermat, Euler, and Gauss – as contemporaries. Hermann Weyl’s writings also influenced me, prompting a desire to explore the relationship between physical and mathematical concepts from a philosophical perspective. Recently, I’ve focused on the history of categories, partly because I’ve contributed to their development and can draw on my memories. This area of mathematics closely aligns with philosophical issues, which I find more compelling than formal logic itself. The central question remains: what guarantees that mathematics conveys truth, and how does it do so consistently121212personal communication? [27, 28].
@@
To do good science, you need that: constant imagination. No prejudice, and as I’ve learned from experience, no fear that your ideas might be silly.

### v2

@@ 1 Hopf algebras, group theory
Cartier had a guiding thread in his work that revolved around groups and Hopf algebras [14, 15]. He viewed Hopf algebras as a powerful tool for understanding various algebraic phenomena. His research often focused on their applications in algebraic topology, representation theory, and mathematical physics. The duality in Hopf algebras, where every algebraic structure has a corresponding co-algebra structure, resonated with Cartier’s broader mathematical philosophy, which emphasized the interconnectedness of different fields of mathematics.
@@ 1.1 On the theory of Formal groups
The essential progress was the introduction of curve modules associated with formal groups and the notion of a typical curve. This novel concept retrospectively explained the arithmetic properties of the classical exponential series and, in particular, the existence of the Artin-Hasse exponential. It also offered a flexible interpretation of the Dieudonné module and allowed me to refine known theorems on Witt groups. The universal typical curve in a Witt group plays a central role in these demonstrations [^8] Personal communication [1].
@@ 1.2 Quantum groups
At that time, I was nearing the end of my mathematical research program and was in a period of waiting. I realized that quantum groups were something entirely new that could be approached with techniques I had previously used (like Lie groups and differential geometry), which I knew well. There was also motivation from physics, even if not always apparent, which reoriented my interests for a good ten years. In fact, I still organize a seminar at the École Polytechnique called ”Quantum Groups and Poisson Geometry,” which continues this work, twenty years later [^9] Personal communication.
@@ 1.3 The Cosmic Galois Groups
The motivic Galois group deals with the automorphisms of certain transcendental numbers, which are similar to the constants appearing in Feynman diagram computations. This observation led me to interpret the group in [12] as a symmetry group governing the fundamental constants of physics. In the Standard Model, these constants are often adjusted empirically, without much mathematical explanation. I believe that this group could express new symmetries among these constants, which might have significant implications for cosmology. My ultimate dream is to unite the ideas in [12] with the motivic Galois group, even if, for now, it’s still an ongoing research program101010personal communication [11].
@@
It took a lot of work and time to obtain a concrete realization of Pierre Cartier’s idea of the Cosmic Galois Group as well as its action on the coupling constants of renormalizable physical theories. The first step, following the introduction of the Hopf algebra of Feynman graphs, was the realization that the renormalization process is, in fact (in the dimensional regularization scheme), identical to the Birkhoff decomposition in pure mathematics, which occurs in the classification of vector bundles on the sphere (see [25]). Once this result was obtained, it again took time, using the Riemann-Hilbert correspondence, to realize ([26]) that there exists a universal Hopf algebra $\mathcal{H}$ behind all these computations111111introducing the category of “equisingular flat connections,” which is shown to be a Tannakian category, meaning it is equivalent to the category of modules over a certain pro-algebraic group. This is the graded dual $\mathcal{H}$ of the universal enveloping algebra of the free graded Lie algebra generated by elements of degree $n$ for each positive $n$. Let $\mathbb{U}$ be the affine group scheme associated with $\mathcal{H}$, and let $\mathcal{H}$ and $\mathbb{U}^{*}=\mathbb{U}\rtimes\mathbb{G}_{m}$ be its semidirect product by the action of the multiplicative group given by the grading. It follows (see [26], Corollary 1.107) that this group acts on the coupling constants of any renormalizable physical theory, thus providing the best model for the Cosmic Galois Group. It gives a concrete realization of the symmetry group of physical theories and has a tantalizing similarity with the motivic Galois group of mixed Tate motives (see [26], Corollary 1.111).
Curiously, one of Cartier’s most cited results is the so-called Cartier theorem, often discussed in the context of algebraic groups over fields of characteristic zero. This theorem, a key result in algebraic geometry and group theory, asserts that algebraic groups over fields of characteristic zero are smooth. It plays an important role in the study of algebraic groups, as smoothness is a fundamental property with significant geometric and arithmetic implications.
In this context, smoothness means that the algebraic group is a smooth variety–its tangent space at every point is well-behaved, which roughly implies that there are no singularities in the variety. This result is foundational in the theory of algebraic groups because smoothness is a desirable property in geometric structures, ensuring that algebraic groups can be analyzed using standard differential-geometric tools (as they behave “well” locally) and that they exhibit favorable properties in representation theory.
The fact that this theorem is only mentioned as a footnote in one of Cartier’s papers [16] is due to its relatively technical nature and the fact that it was not the central focus of his work at the time. Although the proof of smoothness for algebraic groups in characteristic zero was not a primary goal in that paper, it was still an important result for the completeness of his theory, which is why it is briefly noted.
@@ 2 His many key ideas offered to others
What attracts me to the epistemology of mathematics is understanding how mathematics is woven into civilization. I’m interested in how mathematical concepts emerge from societal concerns, reflecting the zeitgeist. One of my students specialized in the history of mathematics in China and worked on her thesis with me and my sinologist brother. André Weil sparked my interest in the history of mathematics, teaching me to view past mathematicians – like Euclid, Archimedes, Fermat, Euler, and Gauss – as contemporaries. Hermann Weyl’s writings also influenced me, prompting a desire to explore the relationship between physical and mathematical concepts from a philosophical perspective. Recently, I’ve focused on the history of categories, partly because I’ve contributed to their development and can draw on my memories. This area of mathematics closely aligns with philosophical issues, which I find more compelling than formal logic itself. The central question remains: what guarantees that mathematics conveys truth, and how does it do so consistently141414personal communication? [28, 29].
@@
To do good science, you need that: constant imagination. No prejudice, and as I’ve learned from experience, no fear that your ideas might be silly [^16] Personal communication.
3 Acknowldgement
We want to warmly thank Mark C. Wilson for his support and suggestions during the preparation of the tribute. We would like to acknowledge the refree’s careful reading and for their suggestion regarding Cartier’s theorem.